Show that f(z)=z2(ez-1) is differentiable at zo=0. Check whether f is analytic at O.


1
Expert's answer
2021-09-07T18:44:13-0400

The derivative of the function "f(z)" is: "f'(z)=2z(e^z-1)+z^2e^z=e^z(z^2+2z)-2z".

"f(0)=0". A necessary conditions for the function to be analytic are: "\frac{\partial f(z,\bar{z}) }{\partial{\bar{z}}}=0". The latter holds. Rewrite the function "f(z)" as: "f(z)=(x+iy)^2(e^{x+iy}-1)=(x^2-y^2+2xiy)(e^xe^{iy}-1)=(x^2-y^2)e^xe^{iy}-(x^2-y^2)+2xiye^xe^{iy}-2xyi=(x^2-y^2)e^x(cos(y)+i\,sin(y))+2xyie^x(cos(y)+i\,sin(y))-2xiy"

In case we split real and imaginary part as: "f(z)=u(x,y)+iv(x,y)" it is obvious that partial derivatives "\frac{\partial u}{\partial x}", "\frac{\partial u}{\partial y}", "\frac{\partial v}{\partial x}", "\frac{\partial v}{\partial y}" are continuous. The latter are sufficient conditions for the function to be analytic. Thus, the function is analytic. In particular, at point 0.